Find the appropriate critical value for a 86% confidence interval for a population mean that is based on 7 observations. Round your answer to three decimal places, e.g., 0.123.
Answer #1

solution

n = 7

Degrees of freedom = df = n - 1 = 7 - 1 = 6

a ) At 86% confidence level the t is ,

= 1 - 86% = 1 - 0.86 = 0.14

  / 2 = 0.14 / 2 = 0.07

t /2,df = t0.07,6 =1.700

answer = 1.700